JamesTDG:

Please leave an explanation for how to do this, I edited the question so I can't cheat with the answer. Triangle Q T S is shown. Angle A T S is a right angle. An altitude is drawn from point T to point R on side Q S to form a right angle. The length of T S is 4 x, the length of Q R is 5, and the length of R S is 11. What is the length of side TS?

KynxKatz:

So wht do you need help wit

ScamTheMan:

Bit confused, any chance you can post a picture of the triangle that's shown? Not sure where R came from if the triangle is QTS

SugarDaddy124:

I cant help you out if I cant see the picture

AZ:

I drew it out based on the information given 

AZ:

But first let's go back to the basics So when we have two right triangles like this there's this formula so for the triangle I'm attaching \(\sf \dfrac{BC}{AC} = \dfrac{AC}{DC}\) and \(\sf \dfrac{BC}{AB} = \dfrac{AB}{DB}\) and \(\sf \dfrac{DC}{AD} = \dfrac{AD}{DB}\)

1 attachment
TypicalFoxGirl:

@AZ I see a "Angle ATS", is that a misspelling or something? I don't see it in your drawing

AZ:

So from my image, the one you would need to use is \(\sf \dfrac{BC}{AB} = \dfrac{AB}{DB}\) So can you apply that formula to your question and solve for x?

AZ:

And remember that the question is asking for the length of TS, so that means you have to plug back x into 4x to get your final answer

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @typicalfoxgirl @AZ I see a "Angle ATS", is that a misspelling or something? I don't see it in your drawing \(\color{#0cbb34}{\text{End of Quote}}\) Good catch. I assumed that said QTS and that's the only thing that would make sense. I forgot to put the right angle for it so here's an updated drawing 

JamesTDG:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ \(\color{#0cbb34}{\text{Originally Posted by}}\) @typicalfoxgirl @AZ I see a "Angle ATS", is that a misspelling or something? I don't see it in your drawing \(\color{#0cbb34}{\text{End of Quote}}\) Good catch. I assumed that said QTS and that's the only thing that would make sense. I forgot to put the right angle for it so here's an updated drawing Created with RaphaëlReply Using Drawing \(\color{#0cbb34}{\text{End of Quote}}\) Thanks for the assistance!

AZ:

You're welcome! Let me know what you get and I can check it for you :)